# Improving robustness to corruptions with multiplicative weight perturbations

### v2

@@ 2.2 Multiplicative weight perturbations simulate input corruptions
Assumption 1.
For each corruption function $\mathbf{g}:\mathcal{X}\rightarrow\mathcal{X}$ in $\mathcal{G}$, there exists a constant $M>0$ such that $||\mathbf{g}(\mathbf{x})-\mathbf{x}||_{2}\leq M$ for all $\mathbf{x}\in\mathcal{X}$.
Assumption 2.
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)}\leq{\color[rgb]{0.99609375,0.26171875,0.234375}{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)},\mathbf{W}^{(h+1)}\right\rangle}_{F}}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}||\mathbf{W}^{(h)}||_{F}^{2}}$ (4)
Here $\otimes$ denotes the outer product of two vectors, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product of two matrices of the same dimension, $||\cdot||_{F}$ is the Frobenius norm, and $\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ is the Jacobian of the per-sample loss with respect to the pre-activation output $\mathbf{z}^{(h)}(\mathbf{x})$ at layer $h$. To prove Lemma 1, we use Assumption 1 and the following assumption about the loss function:
Assumption 3.
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))}\leq{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi}_{\mathbf{g}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (5)
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}(\boldsymbol{\omega};\mathcal{G}(\mathcal{S}))}\leq\mathbb{E}_{\mathbf{g}\sim\mathcal{G}}\left[{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi}_{\mathbf{g}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}}\right]$ (6)
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}_{\boldsymbol{\Xi}}(\boldsymbol{\omega};\mathcal{S})}\vcentcolon=\mathbb{E}_{\boldsymbol{\xi}\sim\boldsymbol{\Xi}}\left[{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}\right]+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (7)
@@
$\min_{\boldsymbol{\omega}}\max_{||\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (8)
@@
$\min_{\boldsymbol{\omega}}\max_{||T_{\boldsymbol{\omega}}^{-1}\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (10)
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}_{\mathrm{max}}(\boldsymbol{\omega};\mathcal{S})}\vcentcolon=\max_{||\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (12)
@@
$\operatorname*{arg\,max}_{||\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}\approx\operatorname*{arg\,max}_{||\boldsymbol{\xi}||_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\left\langle\boldsymbol{\omega}\circ\boldsymbol{\xi},\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\right\rangle$ (13)
@@
$\widehat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\rho\frac{\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{||\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})||_{2}}$ (14)
@@
$\widehat{\boldsymbol{\omega}}=\boldsymbol{\omega}+\boldsymbol{\omega}\circ\widehat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\boldsymbol{\omega}+\rho\frac{\boldsymbol{\omega}^{2}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{||\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})||_{2}}$ (17)
@@ 4 Empirical evaluation
In this section, we assess the corruption robustness of DAMP and ASAM in image classification tasks. We conduct experiments using the CIFAR-10/100 (Krizhevsky, 2009), TinyImageNet (Le and Yang, 2015), and ImageNet (Deng et al., 2009) datasets. For evaluation on corrupted images, we utilize the CIFAR-10/100-C, TinyImageNet-C, and ImageNet-C datasets provided by Hendrycks and Dietterich (2019), as well as ImageNet-$\bar{\text{C}}$ (Mintun et al., 2021), ImageNet-D (Zhang et al., 2024), ImageNet-A (Hendrycks et al., 2021), ImageNet-Sketch (Wang et al., 2019), ImageNet-Drawing (Salvador and Oberman, 2022), and ImageNet-Cartoon (Salvador and Oberman, 2022) datasets, which encapsulate a wide range of corruptions. Detail descriptions of these datasets are provided in Appendix E. We further evaluate the models on adversarial examples generated by the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014). In terms of architectures, we use ResNet18 (He et al., 2016a) for CIFAR-10/100, PreActResNet18 (He et al., 2016b) for TinyImageNet, ResNet50 (He et al., 2016a), ViT-S/16, and ViT-B/16 (Dosovitskiy et al., 2021) for ImageNet. We ran all experiments on a single machine with 8 Nvidia V100 GPUs. Appendix F includes detailed information for each experiment.
@@ 4.3 Benchmark results
Table 1: DAMP surpasses the baselines on corrupted images in most cases and on average. We report the predictive errors (lower is better) averaged over 3 seeds for the ResNet50 / ImageNet experiments. We evaluate the models on IN-C, IN-$\overline{\mathrm{C}}$, IN-A, IN-D, IN-Sketch, IN-Drawing, IN-Cartoon and adversarial examples generated by FGSM. For FGSM, we use $\epsilon=2/224$. For IN-C and IN-$\overline{\mathrm{C}}$, we report the results averaged over all corruption types and severity levels. The Avg column displays the average of all previous columns except IN-Clean. We use 90 epochs and the basic Inception-style preprocessing for all experiments.
Method Error (%) $\downarrow$
IN-Clean FGSM IN-A IN-C IN-$\overline{\mathrm{C}}$ IN-Cartoon IN-D IN-Drawing IN-Sketch Avg
Dropout $23.6_{\pm 0.2}$ $90.7_{\pm 0.2}$ $\mathbf{95.7_{\pm<0.1}}$ $61.7_{\pm 0.2}$ $61.6_{\pm<0.1}$ $49.6_{\pm 0.2}$ $88.9_{\pm<0.1}$ $77.4_{\pm 1.3}$ $78.3_{\pm 0.3}$ 75.5
DAMP $23.8_{\pm<0.1}$ $\mathbf{88.3_{\pm 0.1}}$ $96.2_{\pm<0.1}$ $\mathbf{58.6_{\pm 0.1}}$ $\mathbf{58.7_{\pm<0.1}}$ $\mathbf{44.4_{\pm<0.1}}$ $88.7_{\pm<0.1}$ $\mathbf{71.1_{\pm 0.5}}$ $\mathbf{76.3_{\pm 0.2}}$ 72.8
SAM $23.2_{\pm<0.1}$ $90.4_{\pm 0.2}$ $96.6_{\pm 0.1}$ $60.2_{\pm 0.2}$ $60.7_{\pm 0.1}$ $47.6_{\pm 0.1}$ $\mathbf{88.3_{\pm 0.1}}$ $74.8_{\pm<0.1}$ $77.5_{\pm 0.1}$ 74.5
ASAM $\mathbf{22.8_{\pm 0.1}}$ $89.7_{\pm 0.2}$ $96.8_{\pm 0.1}$ $58.9_{\pm 0.1}$ $59.2_{\pm 0.1}$ $45.5_{\pm<0.1}$ $88.7_{\pm 0.1}$ $72.3_{\pm 0.1}$ $76.4_{\pm 0.2}$ 73.4
@@
Table 2: ViT-S16 / ImageNet (IN) with basic Inception-style data augmentations. Due to the high training cost, we report the predictive error (lower is better) of a single run. We evaluate corruption robustness of the models using IN-A, IN-D, IN-Sketch, IN-Drawing, IN-Cartoon and adversarial examples generated by FGSM. For IN-C and IN-$\overline{\mathrm{C}}$, we report the results averaged over all corruption types and severity levels. For FGSM, we use $\epsilon=2/224$. We also report the runtime of each experiment, showing that SAM and ASAM take twice as long to run than DAMP and AdamW given the same number of epochs. The Avg column displays the average of all previous columns except IN-Clean. DAMP produces the most robust model on average.
Method #Epochs Runtime Error (%) $\downarrow$
IN-Clean FGSM IN-A IN-C IN-$\overline{\mathrm{C}}$ IN-Cartoon IN-D IN-Drawing IN-Sketch Avg
Dropout 100 20.6h 28.55 93.47 93.44 65.87 64.52 50.37 91.15 79.62 88.06 78.31
200 41.1h 28.74 90.95 93.33 66.90 64.83 51.23 92.56 81.24 87.99 78.63
DAMP 100 20.7h 25.50 92.76 92.92 57.85 57.02 44.78 88.79 69.92 83.16 73.40
200 41.1h 23.75 84.33 90.56 55.58 55.58 41.06 87.87 68.36 81.82 70.65
SAM 100 41h 23.91 87.61 93.96 55.56 55.93 42.53 88.23 69.53 81.86 71.90
ASAM 100 41.1h 24.01 85.85 92.99 55.13 55.64 40.74 89.03 67.80 81.47 71.08
@@
Table 3: ViT / ImageNet (IN) with MixUp and RandAugment. We train ViT-S16 and ViT-B16 on ImageNet from scratch with advanced data augmentations (DAs). We evaluate the models on IN-C, IN-$\overline{\mathrm{C}}$, IN-A, IN-D, IN-Sketch, IN-Drawing, IN-Cartoon and adversarial examples generated by FGSM. For FGSM, we use $\epsilon=2/224$. For IN-C and IN-$\overline{\mathrm{C}}$, we report the results averaged over all corruption types and severity levels. The Avg column displays the average of all previous columns except IN-Clean. These results indicate that: (i) DAMP can be combined with modern DA techniques to further enhance robustness; (ii) DAMP is capable of training large models like ViT-B16; (iii) given the same amount of training time, it is better to train a large model (ViT-B16) using DAMP than to train a smaller model (ViT-S16) using SAM/ASAM.
Model Method #Epochs Runtime Error (%) $\downarrow$
IN-Clean FGSM IN-$\overline{\mathrm{C}}$ IN-A IN-C IN-Cartoon IN-D IN-Drawing IN-Sketch Avg
ViT-S16 Dropout 500 111h 20.25 62.45 40.85 84.29 44.72 34.35 86.59 56.31 71.03 60.07
DAMP 500 111h 20.09 59.87 39.30 83.12 43.18 34.01 84.74 54.16 68.03 58.30
SAM 300 123h 20.17 59.92 40.05 83.91 44.04 34.34 85.99 55.63 70.85 59.34
ASAM 300 123h 20.38 59.38 39.44 83.64 43.41 33.82 85.41 54.43 69.13 58.58
ViT-B16 Dropout 275 123h 20.41 56.43 39.14 82.85 43.82 33.13 87.72 56.15 71.36 58.83
DAMP 275 124h 19.36 55.20 37.77 80.49 41.67 31.63 87.06 52.32 67.91 56.76
SAM 150 135h 19.84 61.85 39.09 82.69 43.53 32.95 88.38 55.33 71.22 59.38
ASAM 150 136h 19.40 58.87 37.41 82.21 41.18 30.76 88.03 51.84 69.54 57.48

### v3

@@ 2.2 Multiplicative weight perturbations simulate input corruptions
Assumption 1 (Bounded corruption).
For each corruption function $\mathbf{g}:\mathcal{X}\rightarrow\mathcal{X}$ in $\mathcal{G}$, there exists a constant $M>0$ such that $\|\mathbf{g}(\mathbf{x})-\mathbf{x}\|_{2}\leq M$ for all $\mathbf{x}\in\mathcal{X}$.
Assumption 2 (Transferable robustness).
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)}\leq{\color[rgb]{0.99609375,0.26171875,0.234375}{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)},\mathbf{W}^{(h+1)}\right\rangle}_{F}}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}\|\mathbf{W}^{(h)}\|_{F}^{2}}$ (4)
Here $\otimes$ denotes the outer product of two vectors, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product of two matrices of the same dimension, $\|\cdot\|_{F}$ is the Frobenius norm, and $\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)$ is the Jacobian of the per-sample loss with respect to the pre-activation output $\mathbf{z}^{(h)}(\mathbf{x})$ at layer $h$. To prove Lemma 1, we use Assumption 1 and the following assumption about the loss function:
Assumption 3 (Lipschitz-continuous objective input gradients).
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))}\leq{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi}_{\mathbf{g}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}}{2}\|\boldsymbol{\omega}\|_{F}^{2}}$ (5)
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}(\boldsymbol{\omega};\mathcal{G}(\mathcal{S}))}\leq\mathbb{E}_{\mathbf{g}\sim\mathcal{G}}\left[{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi}_{\mathbf{g}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{C_{\mathbf{g}}}{2}\|\boldsymbol{\omega}\|_{F}^{2}}\right]$ (6)
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}_{\boldsymbol{\Xi}}(\boldsymbol{\omega};\mathcal{S})}\vcentcolon=\mathbb{E}_{\boldsymbol{\xi}\sim\boldsymbol{\Xi}}\left[{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}\right]+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}\|\boldsymbol{\omega}\|_{F}^{2}}$ (7)
@@
$\min_{\boldsymbol{\omega}}\max_{\|\boldsymbol{\xi}\|_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}\|\boldsymbol{\omega}\|_{F}^{2}}$ (8)
@@
$\min_{\boldsymbol{\omega}}\max_{\|T_{\boldsymbol{\omega}}^{-1}\boldsymbol{\xi}\|_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}\|\boldsymbol{\omega}\|_{F}^{2}}$ (10)
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}_{\mathrm{max}}(\boldsymbol{\omega};\mathcal{S})}\vcentcolon=\max_{\|\boldsymbol{\xi}\|_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}\|\boldsymbol{\omega}\|_{F}^{2}}$ (12)
@@
$\operatorname*{arg\,max}_{\|\boldsymbol{\xi}\|_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}\approx\operatorname*{arg\,max}_{\|\boldsymbol{\xi}\|_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\left\langle\boldsymbol{\omega}\circ\boldsymbol{\xi},\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\right\rangle$ (13)
@@
$\widehat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\rho\frac{\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{\|\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\|_{2}}$ (14)
@@
$\widehat{\boldsymbol{\omega}}=\boldsymbol{\omega}+\boldsymbol{\omega}\circ\widehat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\boldsymbol{\omega}+\rho\frac{\boldsymbol{\omega}^{2}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{\|\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\|_{2}}$ (17)
@@ 4 Empirical evaluation
In this section, we assess the corruption robustness of DAMP and ASAM in image classification tasks. We conduct experiments using the CIFAR-10/100 (Krizhevsky, 2009), TinyImageNet (Le and Yang, 2015), and ImageNet (Deng et al., 2009) datasets. For evaluation on corrupted images, we utilize the CIFAR-10/100-C, TinyImageNet-C, and ImageNet-C datasets provided by Hendrycks and Dietterich (2019), as well as ImageNet-$\overline{\text{C}}$ (Mintun et al., 2021), ImageNet-D (Zhang et al., 2024), ImageNet-A (Hendrycks et al., 2021), ImageNet-Sketch (Wang et al., 2019), ImageNet-{Drawing, Cartoon} (Salvador and Oberman, 2022), and ImageNet-Hard (Taesiri et al., 2023) datasets, which encapsulate a wide range of corruptions. Detail descriptions of these datasets are provided in Appendix E. We further evaluate the models on adversarial examples generated by the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014). In terms of architectures, we use ResNet18 (He et al., 2016a) for CIFAR-10/100, PreActResNet18 (He et al., 2016b) for TinyImageNet, ResNet50 (He et al., 2016a), ViT-S/16, and ViT-B/16 (Dosovitskiy et al., 2021) for ImageNet. We ran all experiments on a single machine with 8 Nvidia V100 GPUs. Appendix F includes detailed information for each experiment.
@@ 4.3 Benchmark results
Table 1: DAMP surpasses the baselines on corrupted images in most cases and on average. We report the predictive errors (lower is better) averaged over 3 seeds for the ResNet50 / ImageNet experiments. Subscript numbers represent standard deviations. We evaluate the models on IN-{C, $\overline{\text{C}}$, A, D, Sketch, Drawing, Cartoon, Hard}, and adversarial examples generated by FGSM. For FGSM, we use $\epsilon=2/224$. For IN-{C, $\overline{\text{C}}$}, we report the results averaged over all corruption types and severity levels. We use 90 epochs and the basic Inception-style preprocessing for all experiments.
Method Clean Error (%) $\downarrow$ Corrupted Error (%) $\downarrow$
FGSM A C $\overline{\text{C}}$ Cartoon D Drawing Sketch Hard Avg
Dropout $23.6_{\color[rgb]{.25,.25,.25}0.2}$ $90.7_{\color[rgb]{.25,.25,.25}0.2}$ $\mathbf{95.7_{\color[rgb]{.25,.25,.25}<0.1}}$ $61.7_{\color[rgb]{.25,.25,.25}0.2}$ $61.6_{\color[rgb]{.25,.25,.25}<0.1}$ $49.6_{\color[rgb]{.25,.25,.25}0.2}$ $88.9_{\color[rgb]{.25,.25,.25}<0.1}$ $77.4_{\color[rgb]{.25,.25,.25}1.3}$ $78.3_{\color[rgb]{.25,.25,.25}0.3}$ $85.8_{\color[rgb]{.25,.25,.25}0.1}$ $76.6$
DAMP $23.8_{\color[rgb]{.25,.25,.25}<0.1}$ $\mathbf{88.3_{\color[rgb]{.25,.25,.25}0.1}}$ $96.2_{\color[rgb]{.25,.25,.25}<0.1}$ $\mathbf{58.6_{\color[rgb]{.25,.25,.25}0.1}}$ $\mathbf{58.7_{\color[rgb]{.25,.25,.25}<0.1}}$ $\mathbf{44.4_{\color[rgb]{.25,.25,.25}<0.1}}$ $88.7_{\color[rgb]{.25,.25,.25}<0.1}$ $\mathbf{71.1_{\color[rgb]{.25,.25,.25}0.5}}$ $\mathbf{76.3_{\color[rgb]{.25,.25,.25}0.2}}$ $85.3_{\color[rgb]{.25,.25,.25}0.2}$ $\mathbf{74.2}$
SAM $23.2_{\color[rgb]{.25,.25,.25}<0.1}$ $90.4_{\color[rgb]{.25,.25,.25}0.2}$ $96.6_{\color[rgb]{.25,.25,.25}0.1}$ $60.2_{\color[rgb]{.25,.25,.25}0.2}$ $60.7_{\color[rgb]{.25,.25,.25}0.1}$ $47.6_{\color[rgb]{.25,.25,.25}0.1}$ $\mathbf{88.3_{\color[rgb]{.25,.25,.25}0.1}}$ $74.8_{\color[rgb]{.25,.25,.25}<0.1}$ $77.5_{\color[rgb]{.25,.25,.25}0.1}$ $85.8_{\color[rgb]{.25,.25,.25}0.3}$ $75.8$
ASAM $\mathbf{22.8_{\color[rgb]{.25,.25,.25}0.1}}$ $89.7_{\color[rgb]{.25,.25,.25}0.2}$ $96.8_{\color[rgb]{.25,.25,.25}0.1}$ $58.9_{\color[rgb]{.25,.25,.25}0.1}$ $59.2_{\color[rgb]{.25,.25,.25}0.1}$ $45.5_{\color[rgb]{.25,.25,.25}<0.1}$ $88.7_{\color[rgb]{.25,.25,.25}0.1}$ $72.3_{\color[rgb]{.25,.25,.25}0.1}$ $76.4_{\color[rgb]{.25,.25,.25}0.2}$ $\mathbf{85.2_{\color[rgb]{.25,.25,.25}0.1}}$ $74.7$
@@
Table 2: ViT-S16 / ImageNet (IN) with basic Inception-style data augmentations. Due to the high training cost, we report the predictive error (lower is better) of a single run. We evaluate corruption robustness of the models using IN-{C, $\overline{\text{C}}$, A, D, Sketch, Drawing, Cartoon, Hard}, and adversarial examples generated by FGSM. For IN-{C, $\overline{\text{C}}$}, we report the results averaged over all corruption types and severity levels. For FGSM, we use $\epsilon=2/224$. We also report the runtime of each experiment, showing that SAM and ASAM take twice as long to run than DAMP and AdamW given the same number of epochs. DAMP produces the most robust model on average.
Method #Epochs Runtime Clean Error (%) $\downarrow$ Corrupted Error (%) $\downarrow$
FGSM A C $\overline{\text{C}}$ Cartoon D Drawing Sketch Hard Avg
Dropout 100 20.6h 28.55 93.47 93.44 65.87 64.52 50.37 91.15 79.62 88.06 87.19 79.30
200 41.1h 28.74 90.95 93.33 66.90 64.83 51.23 92.56 81.24 87.99 87.60 79.63
DAMP 100 20.7h 25.50 92.76 92.92 57.85 57.02 44.78 88.79 69.92 83.16 85.65 74.76
200 41.1h 23.75 84.33 90.56 55.58 55.58 41.06 87.87 68.36 81.82 84.18 72.15
SAM 100 41h 23.91 87.61 93.96 55.56 55.93 42.53 88.23 69.53 81.86 85.54 73.42
ASAM 100 41.1h 24.01 85.85 92.99 55.13 55.64 40.74 89.03 67.80 81.47 84.31 72.55
@@
Table 3: ViT / ImageNet (IN) with MixUp and RandAugment. We train ViT-S16 and ViT-B16 on ImageNet from scratch with advanced data augmentations (DAs). We evaluate the models on IN-{C, $\overline{\text{C}}$, A, D, Sketch, Drawing, Cartoon, Hard}, and adversarial examples generated by FGSM. For FGSM, we use $\epsilon=2/224$. For IN-{C, $\overline{\text{C}}$}, we report the results averaged over all corruption types and severity levels. These results indicate that: (i) DAMP can be combined with modern DA techniques to further enhance robustness; (ii) DAMP is capable of training large models like ViT-B16; (iii) given the same amount of training time, it is better to train a large model (ViT-B16) using DAMP than to train a smaller model (ViT-S16) using SAM/ASAM.
Model Method #Epochs Runtime Clean Error (%) $\downarrow$ Corrupted Error (%) $\downarrow$
FGSM $\overline{\text{C}}$ A C Cartoon D Drawing Sketch Hard Avg
ViT S16 Dropout 500 111h 20.25 62.45 40.85 84.29 44.72 34.35 86.59 56.31 71.03 80.87 62.38
DAMP 500 111h 20.09 59.87 39.30 83.12 43.18 34.01 84.74 54.16 68.03 80.05 60.72
SAM 300 123h 20.17 59.92 40.05 83.91 44.04 34.34 85.99 55.63 70.85 80.18 61.66
ASAM 300 123h 20.38 59.38 39.44 83.64 43.41 33.82 85.41 54.43 69.13 80.50 61.02
ViT B16 Dropout 275 123h 20.41 56.43 39.14 82.85 43.82 33.13 87.72 56.15 71.36 79.13 61.08
DAMP 275 124h 19.36 55.20 37.77 80.49 41.67 31.63 87.06 52.32 67.91 78.69 59.19
SAM 150 135h 19.84 61.85 39.09 82.69 43.53 32.95 88.38 55.33 71.22 79.48 61.61
ASAM 150 136h 19.40 58.87 37.41 82.21 41.18 30.76 88.03 51.84 69.54 78.83 59.85